OpenStudy (anonymous):

A minimum energy of 242 kJ/mol is required to break the chlorine–chlorine bond in Cl2. what type of radiation is this

OpenStudy (aaronq):

242 kJ/mol (which is the energy necessary to break one mole of Cl-Cl bonds) so find the energy of a single bond, this is energy of one photon; use planck's equation \(E_{photon}=h*\dfrac{c}{\lambda}\) h= plancks constant = \(6.626*10^{-34}~J*s\) c= the speed of light (in vacuum) =\(3.0*10^8~m/s\) \(\lambda\) =wavelength find the wavelength of the one photon Look up the type of electromagnetic radiation it is